OpenStudy (anonymous):

If g(x) = x2, which expression is equal to g(x + 1)?

OpenStudy (anonymous):

IS IT \(\large g(x)=x^2\)?

OpenStudy (anonymous):

IF SO, THEN \(\huge \color {cyan}{ g(x+1)=(x+1)^2}\).

OpenStudy (anonymous):

Because you're just substituting it.

OpenStudy (anonymous):

thank you!!